OpenStudy (anonymous):

Find the line of symmetry of the parabola defined by 3x^2-6x+8. I will give a medal

OpenStudy (anonymous):

Give me a sec to write it on paper

OpenStudy (anonymous):

y = -3x^2 -6x -8 . The axis of symmetry is defined by x = -b/2a. . -b/2a = -(-6)/(2*(-3) = 6/-6 = -1 . Then substitute x=-6 to find 'y'. . y = -3(-1^2) -(6-1) -8 y = -3 +6 -8 y = -5 . The vertex is: (-1,-5).

OpenStudy (anonymous):

Thank you so much!! @grrrxxiii

OpenStudy (anonymous):

I hope that is correct. Please let me know so I can do some research if I am wrong.

OpenStudy (anonymous):

no what he said is wrong ok so the derivative is 6x-6 which is =0 so x=1 Any y=17

OpenStudy (anonymous):

sorry y=5***

OpenStudy (anonymous):

The answer my teacher gave me was x=1. that is all she said

OpenStudy (anonymous):

Can you show me the steps? I think that is where i am confused @joyraheb

OpenStudy (anonymous):

have you took derivative?

OpenStudy (anonymous):

I dont know what that is

OpenStudy (anonymous):

ok forget so it is defined by x=-2b/a

OpenStudy (anonymous):

i mean -b/2a

OpenStudy (anonymous):

Okay whats next?

OpenStudy (anonymous):

amd a=6 and b=-3 so you get 6/-(-2*3) which is 1

OpenStudy (anonymous):

thats x so then you plug 1 for x in the equ to find y

OpenStudy (anonymous):

Oh ok!! Thank you soooooo much! i still have more if you're available to help?

OpenStudy (anonymous):

so 3-6+8=5 so y=5 so the coord are (1,5)

OpenStudy (anonymous):

ye no prob

OpenStudy (anonymous):

What about what is the factored form of the equation 3x^2+6-10

OpenStudy (anonymous):

you mean 6x-10?

OpenStudy (anonymous):

oops i meant x not 6

OpenStudy (anonymous):

@joyraheb

OpenStudy (anonymous):

ok so use delta=b^2 -4ac

OpenStudy (anonymous):

What is that? idk what delta is